OpenStudy (ashy98):

Evaluate this expression for the values given: Find x - 3y if x = 3 and y = -2 A.) -9 B.) -3 C.) 3 D.) 9

OpenStudy (dangerousjesse):

\(\large 3-3\times -2=x\)

OpenStudy (dangerousjesse):

Be sure to use PEMDAS :)

OpenStudy (ashy98):

9?

OpenStudy (dangerousjesse):

Yes :)

OpenStudy (ashy98):

yay & could you help me with some more??

OpenStudy (dangerousjesse):

Sure

OpenStudy (ashy98):

1.) Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3 A.) -185 D.) -77 C.) -23 D.) 85 2.) Evaluate the following expression using the values given: Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2

OpenStudy (dangerousjesse):

\(\large 3\times 3- 2\times 5^2-3\times 3^3+-3^4\)

OpenStudy (dangerousjesse):

\(\large 5\times 5\) and \(\large 3\times 3\times 3\)

OpenStudy (dangerousjesse):

Then \(\large 3\times 3\), \(\large 2\times 25\), and \(\large 3\times 37\)

OpenStudy (ashy98):

for 1 i keep getting -41

OpenStudy (dangerousjesse):

Then solve \(\large 9-50-111+81\) to get your answer.. No, that's not it :)

OpenStudy (ashy98):

=-71?

OpenStudy (dangerousjesse):

Yes, that's what I got.. But Wolfram said -201, so now I'm not sure -.-

OpenStudy (ashy98):

i hate algebra lol -_-